Operations Report – February 10th

Crew 205 Operations Report 10 – FEB – 2019

SOL: 0

Name of person filing report: Veronica Trivino

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on this evening by 1800 (10thFeb), Stopped by this morning 0800 (10thFeb)

Solar— SOC % – Turned off Gen (10th Feb) 70%, Turned on Gen (10th Feb) 20%

Diesel Reading – 25%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 90%, ~ 495 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 148 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01410586 units

Toilet tank emptied: yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49 hours

Beginning charge: 100%

Ending charge: 88%

Currently charging: Yes

Curiosity rover used: Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 63%

Currently charging: Yes

Spirit rover used: Used

Hours: 73

Beginning charge: 100%

Ending charge: 85%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Driven to Hanksville 4 times to get water

Crew Car used and why, where? Not used

General notes and comments: Hab Car is operational and back from workshop

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Operations Report – February 9th

Crew 204 Operations Report 09 – FEB – 2019

SOL: 12

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: Generator

Notes on non-nominal systems: Generator: Non- nominal. The alternator needs to be replaced and the coolant system flushed and hoses replaced. All parts already arrived.

Generator (hours run): Turned on this evening by 1700 (9th Feb), Stopped by this morning 0830 (9th Feb)

Solar— SOC % – Turned off Gen (9th Feb) 72%, Turned on Gen (9th Feb) 55%

Diesel Reading – 28%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 50%, ~ 250 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 159 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01409959 units

Toilet tank emptied: yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 48.5 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not used

Hours: 72.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Curiosity returned yesterday with a very low charge, compared with Oppy. For long distance EVA., it’s recommended to use Oppy.

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? To town for Dinner

Crew Car used and why, where? To town for Dinner

General notes and comments: HabCar: back on campus. The car was taken to the workshop in order to fix the transmission issue. It was successfully repaired and ready to work. Currently, it doesn’t have a spare tire. The tire is on its way. One of the rear tires had a nail on it, it’s already patched.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Sol Summary – February 10th

Crew 205 Sol Summary Report 10-Jan-2019

Sol: 0

Summary Title: Ready for Mars!

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: Today was a special day: the first day we woke up at the Hab! We spent the night together with Crew 204. The place was crowded but we had a great evening conversing and playing board games together.

We woke up full of energy and ready to prepare the final details for our mission. In the morning, we got our first pictures as a crew and proceeded execute crew handover with Crew 204. We came back to the Hab and settled in to our new home for the next two weeks. We enjoyed blueberry pancakes for breakfast, unpacked our bags, did some deep cleaning, took inventory, and loaded the water tanks.

In the afternoon, Atila and David joined us at the Hab to conduct training on mission protocols and operations. Then we split into two groups for EVA training using the rovers in the field. Finally, we practiced putting on the suits and using the earpieces.

Look Ahead Plan: Tomorrow we enter sim. We will start our new daily routine and plan to conduct our first EVAs. Stay tuned for our future reports. We have two weeks of full of exciting projects coming up! We are ready for Mars!

Anomalies in work: None

Weather: Clear skies all day long, intermittent high intensity winds.

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, EVA request, Operations Report, Greenhab Report, Journalist Report, HSO checklist, Crew Bios and pictures, mission patch

Support Requested: We could not find the astronomy handbook or astronomy laptop (neither in the Hab, Science Dome or RAM)

Sol Summary – February 9th

Crew 204 Sol Summary Report 09- FEB – 2019

Sol: 12

Summary Title: Expanding the family

Author’s name: Kunal Naik (Space_Bot)

Mission Status: Nominal

Sol Activity Summary: The time has come for crew 204 to return home. There are mixed feelings in the mind of the crew members. It was a roller coaster ride but ended quite successfully. The crew woke up a bit later than usual, the last warm Martian breakfast with a sip of Coffee. The first thing on the list for the crew was to clean, clean and clean, the Hab as well as all the operation places. Post cleaning was the wait for the upcoming new crew. We were pretty excited to meet the International Emerging Space Leaders Crew (Crew 205). Finally, the wait was over. At noon our Martian family expanded. After the introductions and enjoying lunch together, our crew got to train them on all the basics to survive and efficiently operate around the Hab. The crew 204 handed over the Hab responsibilities to the new crew. Post training, both crews enjoyed burgers together in Hanksville. The crew 205 is an interdisciplinary team, full of enthusiasm and new exciting projects coming up.

Look Ahead Plan: The crew 204 is scheduled to deorbit around 0800 hours and reach Grand Junction for rendezvous with the spacecraft, moving forward to reenter home sweet home.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA’s location: None

Reports to be file: 1) Operations Report 2) Green Hab Report 3) Commander Report 4) Photos

Support Requested: None

Sol Summary – February 8th

Crew 204 Sol Summary Report 08-FEB -2019

Sol: 11

Summary Title: The End Game

Author’s name: Kunal Naik (Space_Bot)

Mission Status: Nominal

Sol Activity Summary: Time to go home. Is it happiness, eagerness or desperation to see actual beings around us, The Psychological game is going to end and what we learn in 2 weeks? Was it something changing?

Yes, of course not a life changing one but surely a lesson to live on limited water, food supply and strict protocols.

Experiencing some percentage of daily life of Astronauts. The first thing which the crew realized after waking up is the final day of the sim.

The crew had the morning breakfast and were in a rush for their morning EVA, Carbon Flight and MarsBag went for EVA while Cosmos and SpaceBot were in their daily Hab activities. Today’s in charge for cooking was SpaceBot, Post EVA, the crew had lunch together with challenges of Sim as a discussion topic. Today there were 2 EVA’s, one in the morning while the second after lunch. This time, it was SpaceBot and Cosmos turn for the EVA. While other crew members rested for a while at the Hab. Green Positive vibes all over the Green Hab. It’s sad that the crew won’t be able to see the growth of their planted seeds. This is the Ending and probably the last Sol Summary.

Look Ahead Plan: The crew has planned to clean the entire facility from top to bottom, from right to left. The crew will be eagerly waiting to welcome, meet and handing over the HAB to the Crew 205 for future simulation.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA’s location: 1) 516100 E; 4254550 N (Widows Peak) 2) 518250 E;
4255500 N (Quarry Road)

Reports to be file: 1) Operations Report 2) Green Hab Report 3)
Journalist Report 4) EVA Report 5) Photos 6) Mission Summary 7) End
Rotation Food Supply excel sheet

Support Requested: None

Greenhab Report – February 8th

Crew 204 Green Hab Report – 8-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 16.3°C

High: 27.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 171 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 19.4° C / 27 %
2) 12.00 PM 27.0° C / 20 %
3) 15:00 PM 18.0° C / 22 %
4) 18:00 PM 16.3° C / 32 %

Changes to crops: None

Narrative: Out of all the sprouts, fenugreek sprouts have grown faster. Cover shaped leaves are about 5 cm high till today and growing, which make a pretty groundcover between taller, widely spaced crops. and the fact that fenugreek roots are capable of fixing nitrogen in the soil to naturally fertilize the crops around them. All these make the crop a perfect all-rounder.

Harvest: None

Support/supplies needed: None

Operations Report – February 8th

Crew 204 Operations Report 08- FEB – 2019

SOL: 11

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1830 (8th Feb), Stopped by this morning 0900 (8th Feb)

Solar— SOC % – Turned off Gen (8th Feb) 77%, Turned on Gen (8th Feb) 72%

Diesel Reading – 30%

Propane Reading – ~ 55%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 58%, ~ 290 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 168 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01409242 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: 1st EVA

Hours: 48.5 hours

Beginning charge: 100%

Ending charge: 42%

Currently charging: Yes

Curiosity rover used: 1st EVA

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 12%

Currently charging: Yes

Spirit rover used: 2nd EVA

Hours: 72.8

Beginning charge: 100%

Ending charge: 25%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? The car is in Workshop, not at the campus

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Journalist Report – February 8th

Crew 204 Crew Journalist Report 08-FEB-2019

Author: Sonal Baberwal

Title: When we left Earth!

Good Morning Earthlings! The final morning on this soil. Yes, we are returning back to the Earth. Probably the re-entry module has been aligned on the launchpad. But we will miss this place a lot. A completely new world! This isolation acted as a detox from the regular world. It was necessary to have self-analysis. A lot of things have been learnt from this mission.

Another morning starts where the crew had been involved in cleaning the Hab. we had the final two general purpose EVAs. the morning EVA was performed by the Commander and myself followed by the GreenHab Officer and the Crew Engineer.

The day passed by really quickly with no awareness of time! I would like to share a short poem I wrote for MDRS experience, pardon me if not up to date!

“The plants, the lab,
The rovers, the Hab.
The love, the care,
The friendship we share.
The food to cook,
The windows we look.
The signing in and signing off,
The mission support and Cap Com.
The EVAs we risked,
Yes, everything will be missed!”
~ Sonal Baberwal.

Congratulations to the Crew and the mission support for this successful mission. This is Sonal Baberwal, HSO and Crew Journalist reporting live for the last time from the upper deck of the Habitat and signing off on the final day! I wish very good luck and best wishes to Crew 205. Stay connected to hear new stories from Crew 205!

Operations Report- February 7th

Crew 204 Operations Report 07- FEB – 2019

SOL: 10

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1922 (7th Feb),
Stopped by this morning 0630 (7th Feb)

Solar— SOC % – Turned off Gen (7th Feb) Not checked, Turned on Gen
(7th Feb) 77%

Diesel Reading – 32%

Propane Reading – ~ 58%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 60%, ~ 300 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 183 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01408827 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Spirit rover used: Not Used
Hours: 72.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: Not Used
Hours: 47.6 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Not Used
Hours: 72.7 hours

Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? The car is in Workshop

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Greenhan Report- February 7th

Crew 204 Green Hab Report – 7-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.2°C

High: 26.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 183 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 22.2° C / 39 %
2) 12.00 PM 24.9° C / 21 %
3) 15:00 PM 19.9° C / 21 %
4) 18:00 PM 18.2° C / 26 %

Changes to crops: None

Narrative: After almost 12 gallons of water, hours of supplementary light and lots of care later, plants are noted to be in great shape and health.

Harvest: None

Support/supplies needed: None